Coby's crafts has mini wreaths on sale 5 for $12.75. SaveMart is selling them at 6 for $15.60. Who offers the better deal AND by how much
1
Expert's answer
2019-11-24T11:49:40-0500

Firstly, let's find out how much do CobyCraft and SaveMart want for 1 mini wreath.

1) Let's denote the cost of 1 mini wreath by x. Then for the Coby:


5 mini wreaths - $ 12.75

1 mini wreath - $ x


We have a proportion. So let's find "x" .


"x = \dfrac{\$12.75 * 1}{5} =\$ 2.55" . Coby wants $2.55 for 1 mini wreath.

2) For SaveMart proportion is as follows:


6 mini wreaths - $15.60

1 mini wreath - $ x


"x = \dfrac{\$15.60 * 1} {6} = \$2.60" . SaveMart wants $2.60 for 1 mini wreath.


3) "\$ 2.60−\$2.55=\$0.05". Coby's wreaths are cheaper by $0.05 than SaveMarth.

Answer: Coby's wreaths are $2.55 each, SaveMart's wreaths are $2.60, so Coby's are cheaper by 0.05$. This means Coby's deal is better.